Determine the net heat radiated by the black body disc of mass 6g and radius 4cm at 400C, initially kept at 300C.


1
Expert's answer
2021-10-11T16:16:46-0400

Solution;

By Stefan-Boltzmann's law of radiation,the net rate of heat transfer by radiation is;

"\frac {Q_{net}}{t}=\sigma Ae(T_2^4-T_1^4)"

For a black body,e=1

"A=\u03c0r^2=\u03c0\u00d70.04^2="5.026×10-3m2

"\frac{Q_{net}}{t}=5.67×10^{-8}×5.026×10^{-3}×(313^4-303^4)"

"\frac{Q_{net}}{t}=0.33331J/s"